OpenStudy (anonymous):

Simplify: Negative four and one-fifth minus negative thirteen and one-tenth Eight and nine-tenths Negative eight and nine-tenths Negative seventeen and three-tenths Seventeen and three-tenths

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@nuccioreggie

OpenStudy (danjs):

\[-4 \frac{ 1 }{ 5 } - (-13 \frac{ 1 }{ 10 })\]

OpenStudy (danjs):

Turn into improper fractions first \[\frac{ -21 }{ 5 } + \frac{ 131 }{ 10 }\] multiply first fraction by 2/2 \[\frac{ -42 }{ 10 } + \frac{ 131 }{ 10 }\] add \[\frac{ -42 + 131 }{ 10 } = \frac{ 89 }{ 10 } = 8 \frac{ 9 }{ 10 }\]